OpenStudy (anonymous):

18.01 Single Variable Calculus Problem Set 1. Question 1C-4a "Write an equation for the tangent line for the following functions. a) f(x) = 1/(2x+1) at x=1" why isn't the tangent line just f'(1)? (f'(x)= -2/(2x+1)^2), so f'(x)= -2/9 Please explain the reasoning behind the solution "a) f'(1)=-2/9 and f(1)=1/3, so y=-(2/9)(x-1)+1/3=(-2x+5)/9"

OpenStudy (phi):

***why isn't the tangent line just f'(1)? *** the slope of the line is f'(1), which evaluates to -2/9 but there are many lines (an infinite number) (all parallel to themselves) with this slope we want the line that touches the curve at x=1 at x=1, the y value of the point on the curve is y= f(1)= ⅓ so we want the line with slope -2/9 that goes through the point (1,1/3)

OpenStudy (phi):

if you don't know how to find a line given its slope and a point on the line, then you *must* review your algebra (and trig and geometry) before tackling calculus

OpenStudy (anonymous):

f'(x) at x=1 gives a number which is -2/9 . This is the slope of the line we just need. But the question is to find the equation of the line. for finding the equation of a line the requirements are 1.slope 2.A point in (x,y) plane. then the equation of a line in general form can be given as (y-y*) = m(x-x*) where m = slope (x*,y*) is the point on the plane. I hope this helps..